Why the last two is not correct??
Kubin Companys relevant range of per unit are as follows s 21,000 to 25,000 units When it produces and sells 23,000 units, ts average costs Direct materials Direct labor 8.10 $5.10 s 2.66e 6.10 4.68 s 3.6e 2.10 Variable manufacturing Fixed manufacturing Fixed selling expense Fixed administrative expense Sales commissions Variable administrative1.60 1 For 2 l accounting purposes, what is the total amount of product costs incurred to make 23,000 units? For financial accounting purposes, what is the total amount of period costs incurred to sell 23,000 units? 3. For financial 4 For financial accounting purposes, what is the total amount of period costs what is the total amount of product costs incurred to make 25,000 units? ed to sell 21,000 units? 547 500
Answer #1
3
Direct materials 202500 =25000*8.1
Direct labor 127500 =25000*5.1
Variable manufacturing overhead 65000 =25000*2.6
Fixed manufacturing overhead 140300 =23000*6.1
Total amount of product costs 535300
4
Fixed selling expense 105800 =23000*4.6
Fixed administrative expense 82800 =23000*3.6
Sales commissions 44100 =21000*2.1
Variable administrative expense 33600 =21000*1.6
Total amount of period costs 266300